OS AY-2-Moonlght (With Our P Beyonce Radio-Now Playingo x + Take a Test Karissa Moreira Google Chrome htps/www STAT 201-002 Spring 2019 Test.asptestild-1950588848centerwinye Karissa Moreira 1/29/19 6:35 PM Quiz: Quiz Chapter2 Time Remaining: 00:35:43 Submil Quie This Question: 1 p 7 of 11 (7 complete) This Quiz: 11 pts possible A magazine article examined the impact of an applicans ethricity on the likelhood of admission Ethnicity magnet schools programs. Those data are summarized in the Black/Hispanic Admission Decision to a school districts Accepted Waitlisted 141 60 74 453 819 373 1394 a) What percent of all applicants were Asian? The percent of al applicants who were Asian is Round to one decimal place as needed) b)What percent of the students accepted were Asian? Tho percent of students accepted who were Asian is eound to one decimal place as needed ) 디What percent at Alians were accepted? Tho perent of Asam who were occepted isl. Round to one decimal place as needed.) d) whan percentof allshuderts were accepted? The percent of all students who were accepted in Hound to one decimal place as needed ) Chapte 2 Review Chapter 2 Pos Tev PrtSen End PgUp 4 6 7 8 0
Answer #1

Solution:-

a) The percent of all applicant were asians is 32.5%

Total number of applicants = 1394

Total number of Asian applicants = 453

The percent of all applicant were asians = (453/1394)*100 = 32.496%

b) The percent of students accepted were asian is 29.7%.

Total number of accepted students = 202

Total number of asians accepted = 60

The percent of students accepted were asian = (60/202)*100 = 29.703%

c) The percent of asians accepted is 13.3%.

Total number of Asians = 453

Total number of asians accepted = 60

The percent of students accepted were asian = (60/453)*100 = 13.25%

d) The percent of all students who were accepted = 14.5%

Total number of accepted students = 202

Total number of students = 1394

The percent of all students who were accepted = (202/1394)*100 = 14.49%